# Independence Assumption RLS Regularized Least Squares

### v1

@@ C.2 Estimating mutual information from the partial support
$=\frac{1}{Z}\sum_{x\in\tilde{{\mathcal{X}}}}\mu(x)\ln\left(\frac{\mu(x)+\gamma}{\mu^{\otimes}(x)+\gamma}\right)+\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}}}\mu(x)\ln\left(\frac{\mu(x)+\gamma}{\mu^{\otimes}(x)+\gamma}\right)$
@@
$\leq\frac{1}{Z}\sum_{x\in\tilde{{\mathcal{X}}}}\mu(x)\ln\left(\frac{\mu(x)+\gamma}{\mu^{\otimes}(x)+\gamma}\right)+\frac{1}{Z}\,\ln\left(\frac{1+\gamma}{\gamma}\right)\,\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})$
@@
$\leq\frac{1}{Z}\sum_{x\in\tilde{\mathcal{X}}}\left(\mu(x)\ln\left(\frac{\mu(x)}{\mu^{\otimes}(x)}\right)+\gamma\right)+\frac{1}{Z}\,\ln\left(1+\frac{1}{\gamma}\right)\,\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})$ (by Equation 3)
@@
$=\frac{1}{Z}\left(D_{\mathrm{KL}}(\mu,\mu^{\otimes})+|\tilde{\mathcal{X}}|\,\gamma\right)+\frac{1}{Z}\,\ln\left(1+\frac{1}{\gamma}\right)\,\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})~.$
@@
$(ii)\leq\frac{1}{Z}\,\sum_{x\in{\mathcal{X}}^{n}}\xi(x)\,\mu(x)\ln\left(1+\frac{1}{\gamma}\right)=\frac{1-Z}{Z}\,\ln\left(1+\frac{1}{\gamma}\right)~.$
Next, observe that $(iii)\leq\ln(1/Z)$. Finally, putting all together,
$\widehat{I}_{k}(\gamma)$ $\leq\frac{1}{Z}\left(D_{\mathrm{KL}}(\mu,\mu^{\otimes})+|\tilde{\mathcal{X}}|\gamma\right)+\frac{1}{Z}\,\ln\left(1+\frac{1}{\gamma}\right)\left(\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})+1-Z\right)+\ln(1/Z)~.$
@@
$Z\,\widehat{I}_{k}(\gamma)$ $\leq D_{\mathrm{KL}}(\mu,\mu^{\otimes})+|\tilde{\mathcal{X}}|\gamma+\ln\left(1+\frac{1}{\gamma}\right)\left(\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})+1-Z\right)+1-Z$
$\leq D_{\mathrm{KL}}(\mu,\mu^{\otimes})+|\tilde{\mathcal{X}}|\gamma+\ln\left(e+\frac{e}{\gamma}\right)\left(\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})+1-Z\right)$
To complete the proof we need to give a lower bound on $Z$. Note that $Z=1-U_{k}$ by the definition of $Z$ and Proposition˜C.6, and so by Theorem˜C.1
@@
Using this concentration bound together with the choices of $\gamma$ (also setting $\delta_{\mathrm{supp}}=0$ for the first inequality in the main statement) completes the proof of Theorem˜C.4. $\square$
C.4 Expected missing mass under Zipf distribution
@@
Lemma C.7 (Ohannessian and Dahleh, 2010, Theorem 1).
@@
The idea is to use Lemma˜C.7 to give an upper bound on the missing mass. Therefore, we need to establish a lower bound on $\ln F(v)$. For now, abbreviate
@@
We perform an experiment designed to give a data-dependent estimate of the expected missing mass $\mathbb{E}[U_{k}]$ for some specific datasets. Clearly, we cannot simply apply a concentration bound discussed in Section˜C.1 since the complete support of the pseudo joint distribution derived from the LLM is unknown. To this end, we approximate it with a finite support driven by the language model itself. In particular, given a query we sample responses (at temperature 0.90.90.9) until their total probability mass reaches 0.950.950.95 or we reach 100010001000 responses per query. In case of TriviaQA, we performed 123312331233 queries in total. The mean and the median number of unique responses per query was eventually 118.3118.3118.3 and 222222, respectively. In case of the AmbigQA dataset, we performed 700700700 queries, while the mean and the median number of unique responses was 277277277 and 696969, respectively.
@@
which can be computed in practice. In Figure˜7 we present our results in the form of empirical distributions of different quantities, where each observation corresponds to a single query. We compute the bounds for TriviaQA and AmbigQA datasets (see Section˜6 for details about these datasets). From Figure˜7 we can conclude that the expected missing mass for both datasets is very small: Both the missing mass computed on $\tilde{{\mathcal{X}}}$ and the resulting upper bound on $\mathbb{E}[U_{k}]$ are concentrated close to 0, while the cumulative probability of the approximate support $\tilde{{\mathcal{X}}}$ is close to 111 most of the time, showing that our approximations are meaningful.

### v2

@@ E.2 Estimating mutual information from the partial support
$=\frac{1}{Z}\sum_{x\in\tilde{{\mathcal{X}}}}\mu(x)\ln\left(\frac{\mu(x)+\gamma_{1}}{\mu^{\otimes}(x)+\gamma_{1}}\right)+\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}}}\mu(x)\ln\left(\frac{\mu(x)+\gamma_{1}}{\mu^{\otimes}(x)+\gamma_{1}}\right)$
@@
$\leq\frac{1}{Z}\sum_{x\in\tilde{{\mathcal{X}}}}\mu(x)\ln\left(\frac{\mu(x)+\gamma_{1}}{\mu^{\otimes}(x)+\gamma_{1}}\right)+\frac{1}{Z}\,\ln\left(\frac{1+\gamma_{1}}{\gamma_{1}}\right)\,\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})$
@@
$\leq\frac{1}{Z}\sum_{x\in\tilde{\mathcal{X}}}\left(\mu(x)\ln\left(\frac{\mu(x)}{\mu^{\otimes}(x)}\right)+\gamma_{1}\right)+\frac{1}{Z}\,\ln\left(1+\frac{1}{\gamma_{1}}\right)\,\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})$ (by Equation 3)
@@
$=\frac{1}{Z}\left(D_{\mathrm{KL}}(\mu,\mu^{\otimes})+|\tilde{\mathcal{X}}|\,\gamma_{1}\right)+\frac{1}{Z}\,\ln\left(1+\frac{1}{\gamma_{1}}\right)\,\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})~{}.$
@@
$(ii)\leq\frac{1}{Z}\,\sum_{x\in{\mathcal{X}}^{n}}\xi(x)\,\mu(x)\ln\left(1+\frac{1}{\gamma_{1}}\right)=\frac{1-Z}{Z}\,\ln\left(1+\frac{1}{\gamma_{1}}\right)~{}.$
Next, observe that $(iii)\leq\ln(1/Z)$. Finally, term $(iv)$ is controlled through the following fact shown at the end of this section:
Lemma E.7.
Suppose that $\gamma_{1}\geq 0$ while $\gamma_{2}\geq\gamma_{1}+n(1-Z)$. Then,
$\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{\widehat{\mu}^{\otimes}(x)+\gamma_{2}}\right)\leq 0~{}.$
Putting everything together, we obtain
$\widehat{I}_{k}(\gamma_{1},\gamma_{2})$ $\leq\frac{1}{Z}\left(D_{\mathrm{KL}}(\mu,\mu^{\otimes})+|\tilde{\mathcal{X}}|\gamma_{1}\right)+\frac{1}{Z}\,\ln\left(1+\frac{1}{\gamma_{1}}\right)\left(\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})+1-Z\right)+\ln(1/Z)~{}.$
@@
$Z\,\widehat{I}_{k}(\gamma_{1},\gamma_{2})$ $\leq D_{\mathrm{KL}}(\mu,\mu^{\otimes})+|\tilde{\mathcal{X}}|\gamma_{1}+\ln\left(1+\frac{1}{\gamma_{1}}\right)\left(\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})+1-Z\right)+1-Z$
$\leq D_{\mathrm{KL}}(\mu,\mu^{\otimes})+|\tilde{\mathcal{X}}|\gamma_{1}+\ln\left(e+\frac{e}{\gamma_{1}}\right)\left(\mu({\mathcal{X}}^{n}\setminus\tilde{{\mathcal{X}}})+1-Z\right).$
To complete the proof we need to give a lower bound on $Z$. Note that $Z=1-U_{k}$ by the definition of $Z$ and Proposition E.6, and so by Theorem E.1
@@
Using this concentration bound together with the choices of $\gamma$ (also setting $\delta_{\mathrm{supp}}=0$ for the first inequality in the main statement) completes the proof of Theorem E.4. $\square$
Proof of Lemma E.7.
Observe that
$\widehat{\mu}^{\otimes}(x)$ $=(1-\xi(x))\,\prod_{j=1}^{n}\sum_{t\in S:X_{t,j}=x_{j}}\hat{\mu}(X_{t,1},\ldots,x_{j},\ldots,X_{t,n})$
$=\frac{1}{Z^{n}}\,(1-\xi(x))\,\prod_{j=1}^{n}\sum_{t\in S:X_{t,j}=x_{j}}\mu(X_{t,1},\ldots,x_{j},\ldots,X_{t,n})$
$=\frac{1}{Z^{n}}\,(1-\xi(x))\,\prod_{j=1}^{n}\sum_{x^{\prime}\in{\mathcal{X}}^{n}}(1-\xi(x^{\prime}))\,\mathbb{I}\{x^{\prime}_{j}=x_{j}\}\,\mu(x_{1}^{\prime},\ldots,x_{j},\ldots,x_{n}^{\prime})$
$=\frac{1}{Z^{n}}\,(1-\xi(x))\,\prod_{j=1}^{n}\sum_{{x^{\prime}}^{{}^{\backslash j}}}(1-\xi(x^{\prime}_{1},\ldots,x_{j},\ldots,x^{\prime}_{n}))\,\mu(x_{1}^{\prime},\ldots,x_{j},\ldots,x_{n}^{\prime})~{}.$
Now, using that fact that
$\sum_{{x^{\prime}}^{{}^{\backslash j}}}\xi(x^{\prime}_{1},\ldots,x_{j},\ldots,x^{\prime}_{n})\,\mu(x_{1}^{\prime},\ldots,x_{j},\ldots,x_{n}^{\prime})$
$\qquad\leq\sum_{{x^{\prime}}^{{}^{\backslash j}},x_{j}}\xi(x^{\prime}_{1},\ldots,x_{j},\ldots,x^{\prime}_{n})\,\mu(x_{1}^{\prime},\ldots,x_{j},\ldots,x_{n}^{\prime})$
$\qquad=1-Z$
we arrive at
$\widehat{\mu}^{\otimes}(x)$ $\geq\frac{1}{Z^{n}}\,(1-\xi(x))\,\left(\prod_{j=1}^{n}\left(\sum_{{x^{\prime}}^{{}^{\backslash j}}}\mu(x_{1}^{\prime},\ldots,x_{j},\ldots,x_{n}^{\prime})+Z-1\right)\right)_{+}$
$\stackrel{{\scriptstyle(a)}}{{\geq}}\frac{1}{Z^{n}}\,(1-\xi(x))\,\left(\prod_{j=1}^{n}\sum_{{x^{\prime}}^{{}^{\backslash j}}}\mu(x_{1}^{\prime},\ldots,x_{j},\ldots,x_{n}^{\prime})-n(1-Z)\right)_{+}$
$=\frac{1}{Z^{n}}\,(1-\xi(x))\left(\mu^{\otimes}(x)-n\,(1-Z)\right)_{+}$
where to get $(a)$ we used:
Proposition E.8.
For any $p_{1},\ldots,p_{n}\in[0,1]$ and $a\geq 0$, we have
$\prod_{i=1}^{n}(p_{i}-a)\geq\big{(}\prod_{i=1}^{n}p_{i}\big{)}-n\,a~{}.$
Proof.
The statement following by lower-bounding the left-hand side by its linearization in $a$ (derivative at 00), while realizing that it is a convex function of $a$. ∎
The above gives us that
$\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{\widehat{\mu}^{\otimes}(x)+\gamma_{2}}\right)$
$\qquad\leq\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}(1-\xi(x))\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{\frac{1}{Z^{n}}\,(1-\xi(x))\left(\mu^{\otimes}(x)-n(1-Z)\right)_{+}+\gamma_{2}}\right)$
and focusing on the case $11-\xi(x)=$ 1 (otherwise both sides are 00) the above equals to
$\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{\frac{1}{Z^{n}}\,\left(\mu^{\otimes}(x)-n(1-Z)\right)_{+}+\gamma_{2}}\right)$
$\leq\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{(\mu^{\otimes}(x)-n(1-Z))_{+}+\gamma_{2}}\right)$
$\leq\frac{1}{Z}\sum_{x\in{\mathcal{X}}^{n}}\mu(x)\ln\left(\frac{\mu^{\otimes}(x)+\gamma_{1}}{\mu^{\otimes}(x)-n(1-Z)+\gamma_{2}}\right)$
$\leq 0$
by setting $\gamma_{2}\geq\gamma_{1}+n(1-Z)$. ∎
E.4 Expected missing mass under Zipf distribution
@@
Lemma E.9 (Ohannessian and Dahleh, 2010, Theorem 1).
@@
The idea is to use Lemma E.9 to give an upper bound on the missing mass. Therefore, we need to establish a lower bound on $\ln F(v)$. For now, abbreviate
@@
We perform an experiment designed to give a data-dependent estimate of the expected missing mass $\mathbb{E}[U_{k}]$ for some specific datasets. Clearly, we cannot simply apply a concentration bound discussed in Section E.1 since the complete support of the pseudo joint distribution derived from the LLM is unknown. To this end, we approximate it with a finite support driven by the language model itself. In particular, given a query we sample responses (at temperature $0.9$) until their total probability mass reaches $0.95$ or we reach $1000$ responses per query. In case of TriviaQA, we performed $1233$ queries in total. The mean and the median number of unique responses per query was eventually $118.3$ and $22$, respectively. In case of the AmbigQA dataset, we performed $700$ queries, while the mean and the median number of unique responses was $277$ and $69$, respectively.
@@
which can be computed in practice. In Figure 7 we present our results in the form of empirical distributions of different quantities, where each observation corresponds to a single query. We compute the bounds for TriviaQA and AmbigQA datasets (see Section 6 for details about these datasets). From Figure 7 we can conclude that the expected missing mass for both datasets is very small: Both the missing mass computed on $\tilde{{\mathcal{X}}}$ and the resulting upper bound on $\mathbb{E}[U_{k}]$ are concentrated close to 00, while the cumulative probability of the approximate support $\tilde{{\mathcal{X}}}$ is close to $1$ most of the time, showing that our approximations are meaningful.